OpenStudy (anonymous):

Perry can fill a sorter with oranges from the conveyer belt in 10 minutes. While Perry fills the sorter, Gillian takes oranges out of the sorter and puts them in shipping crates. With Gillian taking oranges out of the sorter, it takes 25 minutes for Perry to fill the sorter. Which of the following can be used to determine the amount of time it takes for Gillian to empty the sorter if Perry does not add oranges?

OpenStudy (anonymous):

A. 1/10 - 1/x = 1/25 B. 1/10 + 1/25 = 1/x C. 1/25 - 1/x =1/10 D. 1/10 - 1/x = x/25

OpenStudy (anonymous):

@ganeshie8 @Preetha

OpenStudy (fettiwap1738):

c

OpenStudy (anonymous):

How do you know this?

OpenStudy (fettiwap1738):

i am majik

OpenStudy (anonymous):

Can you give me an explanation?

OpenStudy (fettiwap1738):

nope i am a bit slow

OpenStudy (anonymous):

Alright well I can't really trust an answer without an factual explanation to go alongside it....

OpenStudy (fettiwap1738):

ok but if u gotta prob w/ history i am hella good

OpenStudy (fettiwap1738):

my father is a history teacher

OpenStudy (anonymous):

Im fairly good at history aswell but math screws me up.